OpenStudy (banjoman):

The sum of the ages of Jesse and Kristen is 70 years. 9 years ago, Jesse's age was 3 times Kristen's age. How old is Jesse now? will medal!

OpenStudy (tgstudios):

Think of a number close to 70 divided by 3

OpenStudy (tgstudios):

I already know what it is but think of a few numbers and divide them by 3

OpenStudy (tgstudios):

or take a low number and multiply by 3

OpenStudy (tgstudios):

My only hint to you is that it is not a whole number

OpenStudy (tgstudios):

Are you still here?

OpenStudy (banjoman):

oh yes , I'm just confused as to the process of setting this up right

OpenStudy (tgstudios):

Ok, I always did these kinds of problems by dividing 70 itself by 3 and going off of something close to that

OpenStudy (1davey29):

Let x represent Jesse's age and let y represent Kristen's age. x+y=70. (x-9)=3(y-9), this represents their age's in relation to each other 9 years ago. Solve for y in your second equation, ((x-9)/3)+9=y. Substitute for y, x+((x-9)/3)+9=70. Solve for x, x+((x-9)/3)=61, (3x/3)+((x-9)/3)=61, (3x+x-9)/3=61, 4x-9=183, 4x=192, x=48

OpenStudy (tgstudios):

i was wondering why you were still typing!

OpenStudy (tgstudios):

p.s. banjoman: thats actually not correct

OpenStudy (banjoman):

so , in order to get this right and I apologize for not understanding still, I'm trying to take their combined age which is 70 then divide by three to get close to 9?

OpenStudy (tgstudios):

The actual correct answer, and i used a calculator, is 52.5 for Jesse, and 17.5 for Kristen

OpenStudy (1davey29):

You have to keep in mind that Jesse's age isn't currently three times Kristen's age. It was three times Kristen's age 9 years ago.

OpenStudy (tgstudios):

:( im so sorry i quit

OpenStudy (tgstudios):

oh well, thanks for clarifying me 1davey29

OpenStudy (1davey29):

Np. It's not an uncommon mistake.

OpenStudy (banjoman):

thank you to both of you for your help